OpenStudy (anonymous):

2/5(x+1)=g, Solve for x.

jimthompson5910 (jim_thompson5910):

The first step I would do is multiply both sides by 5 to get 2(x+1) = 5g

jimthompson5910 (jim_thompson5910):

What's next?

OpenStudy (anonymous):

@jim_thompson5910 Hmmmm Divide by 5g??

jimthompson5910 (jim_thompson5910):

The next step is to distribute 2(x+1) = 5g 2x+2 = 5g

jimthompson5910 (jim_thompson5910):

Then what?

OpenStudy (anonymous):

@jim_thompson5910 Oh ok! Oops. Divide by 2?? I think??

jimthompson5910 (jim_thompson5910):

close, you subtract 2 from both sides first

jimthompson5910 (jim_thompson5910):

2(x+1) = 5g 2x+2 = 5g 2x+2-2 = 5g-2 2x = 5g-2

jimthompson5910 (jim_thompson5910):

then you divide both sides by 2

OpenStudy (anonymous):

@jim_thompson5910 Ohhh then you'd get 5g/2 - 1?

jimthompson5910 (jim_thompson5910):

more like \[\Large x = \frac{5g-2}{2}\]

jimthompson5910 (jim_thompson5910):

not sure how you got that

jimthompson5910 (jim_thompson5910):

oh you can break up the fraction and simplify \[\Large x = \frac{5g-2}{2}\] \[\Large x = \frac{5g}{2}-\frac{2}{2}\] \[\Large x = \frac{5g}{2}-1\]

jimthompson5910 (jim_thompson5910):

so I see how you got that now

jimthompson5910 (jim_thompson5910):

so either \[\Large x = \frac{5g-2}{2}\] or \[\Large x = \frac{5g}{2}-1\] work, I prefer the first one, but it doesn't matter

OpenStudy (anonymous):

@jim_thompson5910 Thanks :) I understand. So they both work. Yeah I'd probably prefer the first one now that I think about it for future problems, kinda gets a little confusing. Thanks again!

jimthompson5910 (jim_thompson5910):

which part is confusing you?

OpenStudy (anonymous):

@jim_thompson5910 Nothing its ok! I got it now, just I think it in the future I'll probably do the first answer like you did, and not make it a little more complicated like my answer was

jimthompson5910 (jim_thompson5910):

ah i gotcha

OpenStudy (anonymous):

Thanks again :)

jimthompson5910 (jim_thompson5910):

sure thing